Solve the system of equation for x and y
(3x+y)=7(y-x)
5(3x-y)=x+3

Answer by flame8855(424)

3x+y = 7y-7x
10x = 6y
y = 5/3x
15x-5y = x +3
14x-5y=3
14x - 25/3x=3
x = 9/17
y= 45/51